OpenStudy (anonymous):

The flower garden has the shape of a right triangle. 15ft of perennial border forms the hypotenuse of the triangle, and one leg is 3ft longer than other leg. Find the lengths of the legs. What are the lengths of the legs? ?=ft

OpenStudy (dumbcow):

use pythagorean theorem a^2 + b^2 = c^2 Let x be length of smaller leg x^2 + (x+3)^2 = 15^2 x^2 + (x^2 +6x + 9) = 225 2x^2 +6x - 216 = 0 x^2 + 3x - 108 = 0 (x+12)(x-9) = 0 x = 9

OpenStudy (llort):

H= 15 O = x A = x+3 Darn it, that stupid bovine beat me again

OpenStudy (dumbcow):

haha

OpenStudy (anonymous):

There two answer not just one please

OpenStudy (llort):

the other is 12

OpenStudy (llort):

x+3=12

OpenStudy (anonymous):

got it thanks

OpenStudy (llort):

np